# Neural network prediction of model parameters for strong lensing samples from Hyper Suprime-Cam Survey

### v1

@@ 4.2 Comparison with S19 and S23
After comparing results from the three conventional modelling methods, namely, YL, S19 and S23 with our network, we realise that even though these methods are modelling the same SuGOHI lenses with identical data quality, other than the Einstein radius, a consensus in the predicted parameters is difficult to achieve. We note that this trend is seen, in spite, of other teams having modelled the 10 systems individually and using highly sophisticated methods. A similar inference is made in S23 as well. Such discrepancies could be due to the differences in the actual algorithms involved in modelling the lens systems. For instance, techniques used for modelling the light profile of the lens galaxy and the source, the choice of including or excluding the external shear and using different combinations of the broad-band data. In addition, the actual mass model of the galaxy can be quite complicated than the SIE+external shear model. However, in the absence of the ground truth, one cannot assess which of the methods and their results are more accurate. It may well be that the limitation is inherent to the quality and resolution of the ground-based survey data and better accuracy on the parameters is not possible unless working with data from space based surveys like the Hubble Space Telescope (e.g., Hezaveh et al., 2017).
@@ 5 Summary and Conclusion
In the era of next generation surveys with expected number of strong lenses to be $\mathcal{O}(10^{4})$, we will not have enough resources to model each lens at a time to determine the lens parameters. In this work, we aim to analyse the strong lenses from ground-based imaging surveys in a fast and automated way to prepare ourselves for the big data expected from surveys such as the Rubin LSST. Analysis of images of strong lens systems from a ground-based survey presents a challenge, as the imaging is limited by poor image quality due to atmospheric seeing and low angular resolution. Additionally, increasing depths of the surveys implies increased number of lenses at higher redshifts where the lensed images can often be faint.
@@
To this end, we developed a simple CNN to analyse the strong lenses from the HSC data, a precursor to Rubin LSST, and estimate the lens model parameters. We first trained and tested our network on 60000 HSC-like galaxy-scale simulated lenses to predict the following three parameters of the SIE lens mass model, namely, the Einstein radius, the axis ratio and the position angle of the major axis of the mass distribution. Once we optimise our network on the simulated data, we then tested it on the real galaxy-scale lenses from SuGOHI to predict the aforementioned parameters. We compared our model predictions for 10 SuGOHI lenses that are also modelled by others in the literature with conventional MCMC modelling methods.
Our network performs reasonably well on the simulated lenses in recovering lens model parameters, especially, the Einstein radius which is the most crucial parameter in inferring the mass. For real lenses, we do not have the ground truth for parameters, like we have for the simulated lenses, to compare the predictions from our network in order to gauge its performance. Thus, after applying our network on the sample of 182 SuGOHI (candidate) lenses, we compare our results with the prediction from the YL pipeline. The predicted Einstein radii are generally consistent from the two methods within the errors given by YL. The predictions for the axis ratios and the position angles, however, are not as robust and the degree to which both methods agree varies across the lens sample.
For the 10 SuGOHI lenses which are common to S19, S23 and our analysis with YL and the network, we found similar results as before. The Einstein radius is a fairly robust parameter to infer but the axis ratio and position angle show large variations when inferred even with sophisticated and detailed MCMC modelling carried out on individual lenses (e.g. S19 and S23).
@@
We also note that processing the SuGOHI lenses via YL to remove the central lens light may introduce uncertainties in the lensed images, particularly, for small Einstein radii system, either by leaving a residue or over-subtracting the flux. This can further contribute to the uncertainties in the parameter estimation. In the near future, we plan to work on developing a better method to simultaneously model the lens light along with the lens mass to minimise some of these uncertainties. We also plan to incorporate more of the lens and source parameters in our computation along with analysing how the performance varies as a function of SNR, number of detected images and presence of foreground contaminants. We also look forward to study and implement interpretability tools for CNNs to understand the modelling by our network and its failure modes. We aim to extend this work to upcoming ground based surveys like LSST.
@@ Acknowledgements
We thank Francisco Villaescusa-Navarro, Shreejit Jadhav, Vishal Upendran and Navin Chaurasiya along with Sukanta Bose for useful discussions on the project. PG acknowledges financial support provided by the University Grants Commission (UGC) of India. She is also grateful to IUCAA for providing hospitable environment to students. We acknowledge the use of the high performance computing facility Pegasus at IUCAA for this work. NY and AK thank financial support by Japan Science and Technology Agency AIP Acceleration Research Grant Number JP20317829.
@@
The SuGOHI  [^2] http://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/ lenses and the YattaLens  [^3] https://github.com/astrosonnen/YattaLens pipeline used in this paper are publicly available. The CNN code and simulated images can be made available upon a reasonable request to the corresponding author.
@@ Appendix A Our conventions for Einstein radius and Position Angle
We convert the Einstein radius values quoted in S19 and S23 to our convention (GRAVLENS, Keeton et al., 2000) using the following relation :
$\theta_{\rm E}=\sqrt{\frac{2q}{1+q^{2}}}\theta_{\rm E}^{\rm S19}=\sqrt{\frac{2q}{1+q^{2}}}\theta_{\rm E}^{\rm YL}=\sqrt{\frac{2q}{1+q^{2}}}\frac{2\sqrt{q}}{(1+q)}\theta_{\rm E}^{\rm S23}$ (2)
@@
where, $q$ is the ratio of semi-minor to semi-major axis of the mass distribution of the SIE. We measure the position angle East of North as shown in the Fig. 6 and convert the position angle values quoted in S19 and S23 to this convention before comparing. These conversions were derived by equating the form of the convergence assumed in each of these methods.
Appendix B Analysis of lenses with the lens light
In the Section 2.1, we mentioned the training sample c, where we add simulated arcs (PSF-convolved and Poisson noise added) to the image cutout of the corresponding central HSC galaxy used to model the arcs. Analysing this training sample is important as it resembles the real SuGOHI lenses (not processed by YL), having the real background noise along with the light from the central lens galaxy. When we train our network on this training sample $c$ and test it on the corresponding simulated test sample (see Fig. 7), we find that Einstein radius is not well-constrained. The network recovers the axis ratio and position angle of the mass distribution fairly well because while simulating the lenses, the axis ratio and the position angle of the lens mass model is considered to be the same as that of the lens light. As a result, the presence of the lens light helps the network to estimate the axis ratio and position angle better.
When we test our network on the real lenses available in the SuGOHI database, which inherently contain the lens light, it performs poorly when compared to the predictions from YL (see Fig. 8 and Fig. 9). In the real lenses from the SuGOHI sample, the light coming from the central lens galaxy, either outshines or contaminates the background lensed arcs in most of the cases, which makes it difficult for the network to detect and study the lensed sources. Besides, for real lenses, the axis ratio and position angle of the lens mass model can be quite different from that of the lens light, making the presence of the lens light an issue. This analysis along with Hezaveh et al. (2017), prompted us to perform lens light subtraction and we have used YL for the same.
Figure 7: Performance of our network on a test sample of 2850 lenses where simulated lensed sources are added to the cutouts of central HSC galaxies used as lenses to simulate the lensed arcs. In this case both the training and test set images have similar features represented by the sample image in the first panel. The size of the training and test image cutouts in pixels is $101\times 101$ which corresponds to the angular size of $17$ arcsec $\times$ $17$ arcsec approximately. The second, third and fourth panels describe the predictions by our network for Einstein radius ($\theta_{\rm E}$), axis ratio ($q$) and position angle (PA) respectively as compared to the true values used to generate these simulated images.
Figure 8: Performance of our network on a test set of 182 grade A and B SuGOHI lenses. In this case the training set is the same as described in the Fig. 7. The sample image in the first panel represents the test set consisting of grade A and B lenses available in the SuGOHI database without any further processing. The second, third and fourth panels describe the predictions by our network for Einstein radius ($\theta_{\rm E}$), axis ratio ($q$) and position angle (PA) respectively as compared to the values predicted by YL along with their errors.
Figure 9: Performance of our network on a test set of 25 grade A SuGOHI lenses which is actually a subset of the test set described in the Fig. 8. We can see that the scatter in the predictions of grade A SuGOHI lenses appear to be similar to that of predictions of the combined grade A and B SuGOHI lenses qualitatively and focusing on the grade A lenses does not lead to any significant improvement in the results.

### v2

@@ 4.1 Network performance comparison of 10 real SuGOHI lenses commonly modelled by S19 and S23b
After comparing results from the three conventional modelling methods, namely, YL , S19 and S23b with our network, we reach a similar conclusion to S23b. Even if sophisticated lens modelling methods are used on the same SuGOHI lenses with identical data quality, barring the Einstein radius, there is little consensus amongst the inferred parameters. The parameters of the ellipticity and its position angle, are likely heavily prior and or algorithm dependent. For instance, insufficiencies in modelling the light profile of the lens galaxy and the source, the choice of including or excluding the external shear and using different combinations of the broad-band data, can all heavily affect the answers beyond the quoted statistical uncertainties. In addition, the actual mass model of the galaxy can be quite complicated than the SIE+external shear model. However, in the absence of the ground truth, one cannot assess which of the methods and their results are more accurate. It may well be that the limitation is inherent to the quality and resolution of the ground-based survey data and better accuracy on the parameters is not possible unless working with data from space based surveys like the Hubble Space Telescope (e.g., Hezaveh et al., 2017).
4.2 Comparison of our network and results with studies from the literature
Our work has numerous similarities but also important differences to the growing literature where a neural network is used to analyse parameters of strong lensing systems, particularly from imaging surveys resembling the HSC survey (e.g., S21, S23a, S23b). We present a brief comparison of these methods in this section. The analysis done in P19 focuses on ground based imaging from the Rubin LSST survey, and given that the survey is yet to start, rely on training and testing their network on images where both the lens and the source images are simulated. The studies of S21, S23a, simulate source galaxies and inject them around real lens galaxies from the Subaru HSC survey. The study of S23b applies the network trained and developed in S23a to real SuGOHI lenses.
All the different groups carrying out these studies use simulation pipelines which are independently developed. The simulations in P19 utilize cutouts with size $57$$\times$ $57$ pixels corresponding to 11.4 arcsec on the side, for g, r and i bands of LSST. They use a Sersic profile for both the source and lens galaxies and analyze images for cases including and excluding the lens light. In S21, S23a, S23b, they simulate the strongly lensed systems in g, r i and z bands with a cutout size of $64$ pixels $\times$ $64$ pixels corresponding to 10.8 arcsec on the side, using HSC galaxies as lenses and HUDF galaxies as sources (keeping the lens light in their simulations). We use the SIMCT framework for our analysis and similarly use existing HSC galaxies as lenses but we keep the profiles of our source galaxies parametric, we only use g, r and i bands with a larger cutout size of $101$ pixels $\times$ $101$ pixels corresponding to 17 arcsec on the side, while we exclude the lens light in the majority of our analysis runs separating out the problem of lens light subtraction and lens modeling analysis, although we do test the impact of including lens light in the analysis. However, for predictions of real SuGOHI lenses based on networks trained without lens light, we need to process these images in order to remove the lens light before feeding it to the network.
The network architecture and the training process play a crucial role in determining the accuracy of a neural network. Both P19 and S21 use a simple neural network, while in S23a and S23b they use a ResNet. Along with the differences in the network architectures, all of these studies differ from in the choice of lens model parameters they infer, loss functions, hyperparameters, sizes of the training, validation and test datasets and so on which could potentially lead to differences in their performance. The overall comparison of the results from our analysis using different simulated train samples and their respective test samples with the results from the aforementioned studies is as follows.
P19 use LSST-like simulated data and use a simple neural network to estimate the Einstein radius, position angle and ellipticity of the lenses. When compared to our results on the aforementioned parameters with their LSST gri case, their network seems to show results that are better than the network we present. We believe this excellent performance is related to the use of simulation for both lens and source galaxies considering a Sersic profile for the light. The performances of the networks when lenses include real HSC-lenses superposed on the simulated lensed arcs embedded in real HSC images paints a slightly different picture. The light profiles of real galaxies are often complex and any deviations from the simplicity could be interpreted as a feature of the lenses source introducing systematic biases in the inferred parameters. This also may be part of the reason why the analysis done in P19 on simulated lenses with and without the lens light lead to only a slight degradation of the results. In comparison, our analysis of the SuGOHI lenses, especially for the Einstein radius parameter, with the network trained on the Lenslight sample (see the lower row of Fig. 5) performs worse than the network trained on the HSCempty sample (the lower row of Fig. 4).
In S21, S23a and S23b, they simulate and analyse HSC-like images, similar to our analysis which enables a fairer comparison with these studies than P19. The analysis carried out in S21 using a simple CNN (see fig. 6 in S21) is quite similar to our analysis using the LensLight sample (see the upper row of Fig. 5). Comparing these two plots, we infer that the predictions for the components of ellipticities from both the studies appear qualitatively similar. The use of lens light to determine the ellipticity works best as long as there is little scatter between the ellipticity of the lens light and the lens potential. For the Einstein radius, although S21 results appear slightly better than our results with the LensLight sample, they are comparable with our predictions for the Einstein radius with the HSCempty sample (see the upper row of Fig. 4).
In S23a, the authors build upon the results of S21 with a ResNet which improves the recovery of any spatial offsets between the lens light and mass distribution. However, although the ResNet improved their results on the $x$ and $y$ dramatically (see fig. 4 in S23a) compared to S21, the results on the components of ellipticity remain comparable and also quantitatively similar to our network predictions with the LensLight sample (see the upper row of Fig. 5). The results in S23a, corresponding to the Einstein radius (see fig. 4 in S23a) seem to slightly degrade compared to S21 and also seem to match the performance of the inference of the Einstein radius from our network trained on the LensLight sample. Although S23a additionally include external s, hhe network has difficulty in the inference of the external shear. The performance of the network from S23a with traditional lens modelling methods on grade A SuGOHI sample of lenses was presented in S23b and can be compared with the comparison between our network and traditional lens modelling method (see Appendix  C). We see that the networks in both these studies have comparable performance in reproducing the results from traditional modeling methods for both the Einstein radius and the two components of the ellipticity.
Even though the studies compared in this section, all aim towards a common scientific goal, differences in the simulation methods, datasets, network architectures, lens model parameters and overall results make these studies interesting in their own right to build a consensus on the direction of required efforts. The discrepancy found in the results of modelling of lenses with traditional methods (see Fig. 9) invites more efforts and independent investigations along these lines.
@@ 5 Summary and Conclusion
The next generation imaging surveys are expected to increase the number of known strong lensing systems to $\mathcal{O}(10^{5})$, which will present formidable challenges for modelling each lens system individually. Analysis of images of strong lens systems from a ground-based survey presents a challenge, as the imaging is limited by poor image quality due to atmospheric seeing and low angular resolution. Additionally, increasing depths of the surveys implies increased number of lenses at higher redshifts where the lensed images can often be faint. In this work, we systematically investigate the ability of neural networks to analyse the strong lenses from ground-based imaging surveys in a fast and automated way and its performance depending upon the realism of the training samples used.
@@
We developed a simple CNN to analyse the strong lenses from the HSC data, a precursor to Rubin LSST, and estimate the parameters of the SIE lens mass model, namely, the Einstein radius, the axis ratio and the position angle of the major axis of the mass distribution. We first trained and tested our network on 60000 HSC-like galaxy-scale simulated lenses. We prepared four different training samples with increasing degree of realism: PureSims sample included only the arcs with Poisson noise, the GauNoise sample further included background Gaussian noise, the HSCempty sample included the lensed features on top of randomly selected cutouts from HSC with central “empty" regions, and the LensLight sample also included the light from the lens galaxy.
We trained our network on each of the four types of samples and checked its performance on both the corresponding simulated test samples and the real sample of 182 galaxy-scale SuGOHI lenses (grade A and B) to predict the lens parameters. The performance of the network on SuGOHI lenses was tested against the parameters predicted by YL for these systems, which were analysed for the first time in this paper.
We summarise the results of our the performance of the CNN that we developed and comparison with the literature studies below.
  • Across the investigations carried out in this work, we find that the Einstein radius of a lens system is the most robust parameter that can be inferred, both from traditional modelling as well as machine learning.
  • The performance of the CNN that we have constructed varies from parameter to parameter depending upon the training sample used to train the network. We have shown that the network trained on the HSCempty simulated sample of lenses is the most robust. This network can predict the Einstein radius with a scatter of 10-20 percent, a bias less than 5 percent and an outlier fraction around 10 percent.
  • On the other hand, the network trained on the Lenslight simulated sample predicts ellipticity parameters more accurately. We show, however, that the improved accuracy of the network trained on the Lenslight sample is correlated to the lens light parameters more than the lens mass parameters. If the real lens galaxies have larger misalignment between the lens mass and light parameters, then such networks are likely to have increased scatter. Therefore, for modelling inferences which are independent of the lens light, our results suggest that it is ideal for networks to be trained after lens light subtraction and that accurate lens light subtraction techniques be further investigated.
  • We compared our work with similar studies in the literature, where a neural network is trained on HSC-like data to predict the lens mass parameters. In spite of using completely independent methodology for simulating lenses, differences in choices or assumptions when generating the lens populations, different network architectures and different means of estimating true (or reference) parameters for real lenses, our accuracies and scatter on the predicted lens mass parameters, especially the Einstein radius, are comparable.
  • We have modelled the 182 SuGOHI (Grade A and B) lenses together for the first time using YL , and presented the parameters and their errors. When the network predictions are compared to those from YL , we find that the network trained on HSCempty sample performs the best amongst the alternatives we explore.
  • We further compared the predictions from our network, trained on the HSCempty sample, for 10 SuGOHI lenses that are also modelled by others (S19, S23b) and by us with YL using conventional MCMC modelling methods. We find that the traditional methods agree on the Einstein radius but the two components of ellipticity show larger scatter even when the lenses are modelled on a case-by-case basis using sophisticated modelling techniques (e.g. S19 and S23b). This may either due to the limitation of ground-based data quality, lens model degeneracies or lack of understanding of misalignment between mass and light components.
  • Networks trained on all samples perform well on the simulated samples and somewhat poorly on the real samples. This holds true even in those networks where the simulated lenses include some misalignment between lens mass and light parameters (see results of S23b, which were found to be consistent with our results on the grade A SuGOHI sample) suggesting a) scope for improvement in the quality and diversity of simulated samples and b) possibly, need for development of a more robust network
Overall our network, once trained, takes around a millisecond to predict the lens mass parameters on a single lens-light subtracted image. This makes our analysis equipped for analysing a plethora of strong lenses $\mathcal{O}(10^{5})$, expected from the next generation surveys like LSST in a reasonable amount of computational cost.
In spite of the ongoing progress in large-scale modelling of lenses, some challenges remain to be addressed hitherto. For instance, we note that processing the SuGOHI lenses via YL , to remove the central lens light, may introduce artifacts in the lensed images either by leaving a residue or over-subtracting the flux. Such artifacts can contribute to increased parameter uncertainties, particularly, for small Einstein radius systems. The absence of true model parameters for the real lenses, and the non-consensus of parameters from the conventional modelling techniques make it harder to objectively quantify the network performance.
@@
In future, we plan to incorporate prediction of the additional lens and source parameters along with analysing how the network performance varies as a function of SNR, number of detected images and presence of foreground contaminants. Furthermore, we also are exploring the using the interpretability tools for CNNs to understand the model predictions by our network and its failure modes. We aim to extend this work to upcoming ground based surveys like LSST.
@@ Acknowledgements
We thank Stefan Schuldt, Francisco Villaescusa-Navarro, Vibhore Negi, Shreejit Jadhav, Vishal Upendran and Navin Chaurasiya along with Sukanta Bose for useful discussions on the project. We also thank Stefan Schuldt for sharing the data. PG acknowledges financial support provided by the University Grants Commission (UGC) of India. We acknowledge the use of the high performance computing facility Pegasus at IUCAA for this work. NY and AK thank financial support by Japan Science and Technology Agency AIP Acceleration Research Grant Number JP20317829.
@@
The SuGOHI  [^4] http://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/ lenses and the YL  [^5] https://github.com/astrosonnen/YattaLens pipeline used in this paper are publicly available. The CNN code and simulated images can be made available upon a reasonable request to the corresponding author.
@@ Appendix A Our conventions for Einstein radius and Position Angle
We convert the Einstein radius values quoted in S19 and S23b to our convention (GRAVLENS, Keeton et al., 2000) using the following relation :
$\theta_{\rm E}=\sqrt{\frac{2q}{1+q^{2}}}\theta_{\rm E}^{\rm S19}=\sqrt{\frac{2q}{1+q^{2}}}\theta_{\rm E}^{\rm YL}=\sqrt{\frac{2q}{1+q^{2}}}\frac{2\sqrt{q}}{(1+q)}\theta_{\rm E}^{\rm S23b}$ (2)
@@
where, $q$ is the ratio of semi-minor to semi-major axis of the mass distribution of the SIE. We measure the position angle East of North as shown in the Fig. 9 and convert the position angle values quoted in S19 and S23b to this convention before comparing. These conversions were derived by equating the form of the convergence assumed in each of these methods. We further obtain the two components of ellipticity using the axis ratio ($q$) and the position angle ($\theta$) in the following relations :
$e_{x}=\frac{1-q^{2}}{1+q^{2}}\cos(2\theta)$ (3)
$e_{y}=\frac{1-q^{2}}{1+q^{2}}\sin(2\theta)$ (4)
Appendix B Errors on the network predictions
We quantify the error in our network predictions based on the performance of the network on the simulated test sample shown in Fig. 4. For example, we present a scatter plot of the ratio of the true Einstein radius to the network-predicted value as a function of the latter in Fig. 10. We divide the sample in $10$ bins and calculate the standard deviation of the afore-mentioned ratio in each bin (green solid curve). For a given SuGOHI grade A lens, the network prediction is then multiplied by this standard deviation, in the corresponding bin, to obtain the error on the network-predicted Einstein radius in Fig. 9. We follow a similar procedure for assigning errors on $e_{x}$ and $e_{y}$, except by using the scatter in the difference of the true value and the network value as shown in Fig. 10. The computed standard deviation for $e_{x}$ and $e_{y}$ is used as an error from the network in the corresponding bin in tFig. 9.
Figure 10: For Einstein radius ($\theta_{\rm E}$), the red points represent a scatter of the ratio of the true value and the network predicted value versus the network value, we divide the parameter range shown in $10$ bins and calculate the standard deviation in each bin as shown by the green points. For the ellipticity components ($e_{x}$ and $e_{y}$), the red points represent a scatter of the difference of the true value and the network value versus the network value, we divide the parameter range in $10$ bins similarly and calculate the standard deviation in each bin as shown by the green points.
Appendix C Performance of our LensLight sample on the grade A SuGOHI lenses
We show the performance of our LensLight sample on 182 SuGOHI (grade A and B) lenses in the lower panel of the Fig. 5. In S23b, they train their ResNet on a sample qualitatively similar to our LensLight sample and test it on a sample of 31 grade A SuGOHI lenses. In order to compare with S23b, here we show the performance of our LensLight sample on a sample 25 grade A suGOHI lenses. Our grade A sample is smaller than the S23b sample, since YL does not identify a few of these 31 grade A lenses as lenses and we do not have the parameters from YL to compare our results with for those lenses. If we compare the results shown in the Fig. 11 (we compare our network predictions with YL ) with the fig.4 in S23b (they compare their network predictions with GLEE & GLAD), we can see that our results are comparable with theirs for the Einstein radius and the components of the ellipticity.
Figure 11: Performance of our network (trained with the LensLight sample) on a sample of 25 SuGOHI grade A lenses (not processed by YL ). The range of the parameters is kept similar to the fig.4 of S23b for ease of comparison.
Appendix D Lens model parameters for SuGOHI lenses
The performance on 182 SuGOHI (grade A and B) lenses of our network trained on the HSCempty sample is shown in the lower panel of the Fig. 4. Here, we present the lens model parameters and errors for these SuGOHI lenses as estimated by YL and our network (trained on the HSCempty sample). We only show 25 grade A lenses in the Table 2. The lens model parameters for 157 grade B lenses, along with their errors are provided in the supplementary material.
Table 2: Lens model parameters predicted by our network (CNN) and by YL , along with errors, for 182 SuGOHI grade A and B lenses (we only show grade A lenses here and grade B lenses are provided in the supplementary material.)
Name Method $\theta_{\rm E}$ (arcsec) $e_{x}$ $e_{y}$ HSCJ015618-010747 CNN $0.80\pm 0.11$ $-0.65\pm 0.49$ $-0.44\pm 0.23$ YL $0.87\pm 0.18$ $-0.34\pm 0.26$ $-0.35\pm 0.23$ HSCJ020141-030946 CNN $1.56\pm 0.50$ $0.29\pm 0.19$ $0.02\pm 0.12$ YL $1.49\pm 0.12$ $0.27\pm 0.14$ $0.18\pm 0.15$ HSCJ020241-064611 CNN $1.38\pm 0.34$ $0.08\pm 0.12$ $0.14\pm 0.12$ YL $1.06\pm 0.42$ $-0.20\pm 0.21$ $-0.20\pm 0.19$ HSCJ020955-024442 CNN $0.80\pm 0.10$ $-0.20\pm 0.18$ $0.47\pm 0.25$ YL $0.93\pm 0.17$ $0.06\pm 0.24$ $-0.09\pm 0.21$ HSCJ021737-051329 CNN $0.79\pm 0.10$ $-0.76\pm 0.49$ $-0.11\pm 0.12$ YL $1.17\pm 0.11$ $-0.45\pm 0.09$ $-0.14\pm 0.15$ HSCJ022346-053418 CNN $1.20\pm 0.24$ $0.34\pm 0.27$ $0.45\pm 0.25$ YL $0.98\pm 0.29$ $-0.54\pm 0.25$ $0.08\pm 0.26$ HSCJ022610-042011 CNN $1.12\pm 0.22$ $0.55\pm 0.38$ $0.02\pm 0.12$ YL $1.10\pm 0.17$ $-0.08\pm 0.21$ $-0.02\pm 0.21$ HSCJ023217-021703 CNN $1.03\pm 0.18$ $0.09\pm 0.12$ $0.68\pm 0.41$ YL $1.35\pm 0.35$ $-0.01\pm 0.23$ $-0.42\pm 0.23$ HSCJ023322-020530 CNN $0.97\pm 0.17$ $0.41\pm 0.27$ $0.00\pm 0.12$ YL $1.56\pm 0.08$ $0.01\pm 0.10$ $-0.01\pm 0.09$ HSCJ085046+003905 CNN $1.15\pm 0.23$ $0.47\pm 0.27$ $0.72\pm 0.41$ YL $1.57\pm 0.08$ $0.57\pm 0.08$ $-0.09\pm 0.07$ HSCJ085855-010208 CNN $1.08\pm 0.19$ $-0.11\pm 0.12$ $0.01\pm 0.12$ YL $1.06\pm 0.04$ $-0.03\pm 0.07$ $0.03\pm 0.06$ HSCJ090429-010227 CNN $0.79\pm 0.10$ $0.25\pm 0.19$ $-0.11\pm 0.12$ YL $0.40\pm 0.13$ $0.16\pm 0.33$ $-0.82\pm 0.24$ HSCJ121052-011905 CNN $1.34\pm 0.33$ $0.26\pm 0.19$ $-0.07\pm 0.12$ YL $1.23\pm 0.11$ $0.21\pm 0.17$ $-0.38\pm 0.15$ HSCJ124320-004517 CNN $1.61\pm 0.52$ $-0.21\pm 0.18$ $-0.17\pm 0.17$ YL $1.51\pm 0.23$ $-0.27\pm 0.18$ $0.04\pm 0.16$ HSCJ125254+004356 CNN $1.46\pm 0.36$ $0.31\pm 0.19$ $-0.26\pm 0.17$ YL $1.13\pm 0.11$ $-0.57\pm 0.17$ $0.03\pm 0.25$ HSCJ135138+002840 CNN $2.41\pm 0.32$ $0.31\pm 0.19$ $-0.03\pm 0.12$ YL $2.12\pm 0.06$ $0.11\pm 0.11$ $0.23\pm 0.20$ HSCJ141136-010215 CNN $0.99\pm 0.17$ $-0.29\pm 0.18$ $-0.00\pm 0.12$ YL $1.00\pm 0.05$ $-0.06\pm 0.10$ $-0.49\pm 0.15$ HSCJ142720+001916 CNN $1.36\pm 0.33$ $0.10\pm 0.12$ $0.36\pm 0.25$ YL $1.36\pm 0.13$ $-0.04\pm 0.17$ $0.13\pm 0.17$ HSCJ144320-012537 CNN $0.92\pm 0.16$ $0.28\pm 0.19$ $0.59\pm 0.34$ YL $1.09\pm 0.06$ $-0.19\pm 0.10$ $-0.59\pm 0.13$ HSCJ145242+425732 CNN $0.98\pm 0.17$ $0.25\pm 0.19$ $-0.62\pm 0.29$ YL $1.85\pm 0.45$ $-0.25\pm 0.25$ $-0.02\pm 0.26$ HSCJ223733+005015 CNN $1.51\pm 0.48$ $-0.14\pm 0.12$ $-0.08\pm 0.12$ YL $1.32\pm 0.23$ $0.08\pm 0.18$ $0.15\pm 0.17$ HSCJ230335+003703 CNN $0.83\pm 0.11$ $-0.44\pm 0.24$ $-0.64\pm 0.42$ YL $0.99\pm 0.27$ $-0.50\pm 0.28$ $-0.02\pm 0.21$ HSCJ230521-000211 CNN $1.06\pm 0.19$ $0.20\pm 0.19$ $-0.61\pm 0.29$ YL $1.68\pm 0.02$ $-0.34\pm 0.06$ $-0.03\pm 0.02$ HSCJ233130+003733 CNN $1.35\pm 0.33$ $0.15\pm 0.12$ $-0.15\pm 0.12$ YL $1.43\pm 0.06$ $0.30\pm 0.07$ $-0.15\pm 0.05$ HSCJ233146+013845 CNN $1.06\pm 0.18$ $-0.80\pm 0.49$ $-0.19\pm 0.17$ YL $1.43\pm 0.17$ $-0.43\pm 0.12$ $0.03\pm 0.13$